OpenStudy (anonymous):

for the following points (-15,12) and (-12,10). A. calculate the slope, in the lowest terms, of the line passing through these two points. B. use point slope to find the equation of the line passing through these two points. C. state the equation in point slope form. D. state the equation in slope-intercept form. E. state the x and y intercepts F. graph the equation of the line.

OpenStudy (swissgirl):

right underneath the textbook ull see a draw button in blue click on that

OpenStudy (lgbasallote):

textbook?

OpenStudy (swissgirl):

textbox*

OpenStudy (swissgirl):

ohhh so then you wld attach the file

OpenStudy (anonymous):



OpenStudy (anonymous):

my graph sucks sorry guys

OpenStudy (nubeer):

ok can do few parts for u. (-15,12) (-12,10) A) slope = (y2-y1)/(x2-x1) y2 = 10 , y1 = 12 ,x1 = -15 ,x2 = -12 just plug in the values in formaula u will get the slope.. can u do this one?

OpenStudy (anonymous):

yea ill try.

OpenStudy (nubeer):

ok good.. for B) u have slope and a point..(pick any point given in your question) i will pick (-12, 10) y-y1 = m (x-x1) m = slope , y1 = 20 , x1 = -12... just plug in values u will get the equation